OpenStudy (anonymous):

x squared = 5x minus 6 solve for x solve and explain please

OpenStudy (anonymous):

x^2=5x-6 x^2-5x+6=0 (x-2)(x-3)=0 x=2 x=3

OpenStudy (anonymous):

x^2 = 5x - 6 move all variables so equation = 0 undo the foil and (x-2)(x-3) =0

OpenStudy (anonymous):

so x=2 or x =3 both real answers

OpenStudy (anonymous):

mistype on 6